Direct Labor Variances

The following data relate to labor cost for production of 4,600 cellular telephones:

Actual: 3,090 hrs. at $16.10
Standard: 3,040 hrs. at $16.40

a. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Rate variance $ Favorable
Time variance $ Unfavorable
Total direct labor cost variance $ Favorable

b. The employees may have been less-experienced workers who were paid less than more-experienced workers or poorly trained, thereby resulting in a lower  labor rate than planned. The lower level of experience or training may have resulted in less  efficient performance. Thus, the actual time required was more  than standard.
Answer #1

a) Direct labor rate variance = (16.40-16.10)*3090 = 927 Favorable

Time variance = (3040-3090)*16.4 = 820 Unfavorable

Total direct labor cost variance = (3040*16.40)-(3090*16.10) = 107 Favorable